OpenStudy (atianaaaa):

Use the Remainder theorem to find the remainder when P(X)=x^4-9x^3-5x^2-3x+4 is divided by x+3

OpenStudy (anonymous):

I wrote this out here F(-3)=(-3)^4 -9*(-3^3)-5*(-3^2)-(3*-3) +4 = 81 + 243 - 45 + 9 + 4 = 292 Remainder =292 Does that help you?